∠DEF and ∠ABC are supplementary angles and ∠DEF is four times as large as ∠ABC. Determine the measure of each angle. For some reason I have major problems with these lines. Any help, please.
Found 2 solutions by rfer, stanbon:
Answer by rfer(16322)

x+4x=180
5x=180
x=36
4x=144

Answer by stanbon(75887)

∠DEF and ∠ABC are supplementary angles and ∠DEF is four times as large as ∠ABC. Determine the measure of each angle
-----
Equations:
E + B = 180 degrees
E = 4B
---
Substitute for "E" and solve for "B";
4B + B = 180
5B = 180
ABC = 36 degrees
---
Solve for E:
DEF = 180-36
DEF = 144 degrees
------
Cheers,
Stan H.